# Context-Alignment: Activating and Enhancing LLM Capabilities in Time Series

### v1

@@ C.2 Short-term Forecasting full results
We present the comprehensive short-term forecasting results in Table 11. Under different frequency settings, DECA consistently outperforms most baseline models.
@@
width = 1colspec = Q[0.2em,c]Q[2em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c], colsep = 0.1pt, rowsep = 1.4pt, cells = c, cell11 = c=20.09, cell2,5,8,11,141 = r=3, hline1 = 1-140.15em, hline2,5,8,11,14 = 1-140.05em, hline17 = 1-140.15em, vline3,4,5,6,7,8,9,10,11,12,13,14 = 3,6,9,12,15, vline3,4,5,6,7,8,9,10,11,12,13,14 = 2,5,8,11,14abovepos = -1.2, vline3,4,5,6,7,8,9,10,11,12,13,14 = 4,7,10,13,16belowpos = -1.2, vline3,4,5,6,7,8,9,10,11,12,13,14 = 1abovepos = -1.2, belowpos = -1.2, Methods & DECA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST N-HiTS N-BEATS TimesNet FEDformer Stationary
@@ C.3 Few-shot Forecasting full results
Table 12 presents the detailed results of the few-shot forecasting trained on 5% data across different prediction lengths. Except for DLinear’s strong performance on the ETTm1 dataset, DECA demonstrates significant enhancements in most scenarios across various datasets, outperforming the second-best model, S2IP-LLM, by 6.7%.
@@
colspec = Q[1.5em,c]Q[2.5em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c], colsep = 2.2pt, rowsep = 1.4pt, cells = c, cell11,3,5,7,9,11,13,15,17,19,21,23 = c=2, cell21 = c=2, cell3,8,13,181 = r=5, hline1 = 1-240.15em, hline2,3,8,13,18 = 1-240.05em, hline23 = 1-240.15em, vline2,3,5,7,9,11,13,15,17,19,21,23 = 4-6,9-11,14-16,19-21, vline2,3,5,7,9,11,13,15,17,19,21,23 = 3,8,13,18abovepos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 7,12,17,22belowpos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 1,2abovepos = -1.2, belowpos = -1.2, Methods & DECA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer
@@
Table 13 presents the detailed results of the few-shot forecasting task using 10% of the training data. DECA demonstrates superior performance, outperforming the baseline across nearly all settings, with an average MSE reduction of 7.8% over the next best model, GPT4TS.
@@
colspec = Q[1.5em,c]Q[2.5em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c], colsep = 2.2pt, rowsep = 1.4pt, cells = c, cell11,3,5,7,9,11,13,15,17,19,21,23 = c=2, cell21 = c=2, cell3,8,13,181 = r=5, hline1 = 1-240.15em, hline2,3,8,13,18 = 1-240.05em, hline23 = 1-240.15em, vline2,3,5,7,9,11,13,15,17,19,21,23 = 4-6,9-11,14-16,19-21, vline2,3,5,7,9,11,13,15,17,19,21,23 = 3,8,13,18abovepos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 7,12,17,22belowpos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 1,2abovepos = -1.2, belowpos = -1.2, Methods & DECA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer
@@ C.4 Zero-shot Forecasting full results
Table 14 provides detailed results of the zero-shot forecasting task across different prediction lengths. DECA consistently achieves optimal performance across all settings, with a significant margin. Compared to the second-best method, PatchTST, it shows an average improvement of 13.3%. Additionally, DECA achieves MSE reductions of 18.3%, 17.7%, and 24.3% over other LLM-based methods S2IP-LLM, Time-LLM, and GPT4TS, respectively. This improvement is attributed to DECA’s successful activation of LLMs’ structured understanding of TS data, while the design of the Dual-Scale Context-Alignment GNNs ensures that LLMs grasp the logical relationships in demonstration examples prompt.
@@
colsep = 2.2pt, rowsep = 1.4pt, cells = c, colspec = Q[5em,c]Q[1.5em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c], cell11,3,5,7,9,11,13,15,17 = c=2, cell21 = c=2, cell3,8,13,18,23,28,33,381 = r=5, hline1 = 1-240.15em, hline2,3,8,13,18,23,28,33,38 = 1-240.05em, hline43 = 1-240.15em, vline2,3,5,7,9,11,13,15,17 = 4-6,9-11,14-16,19-21,24-26,29-31,34-36,39-41, vline2,3,5,7,9,11,13,15,17 = 3,8,13,18,23,28,33,38abovepos = -1.2, vline2,3,5,7,9,11,13,15,17 = 7,12,17,22,27,32,37,42belowpos = -1.2, vline2,3,5,7,9,11,13,15,17 = 1,2abovepos = -1.2, belowpos = -1.2, Methods & DECA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet
@@ C.5 Classification Full Results
Table 15 presents the results on 10 multivariate UEA classification datasets. In the time series classification task, we have augmented the baselines with the following methods: XGBoost Chen and Guestrin [2016], Rocket Dempster et al. [2020], LSTNet Lai et al. [2018b], LSSL Gu et al. [2021], LightTS Zhang et al. [2022], Pyraformer Liu et al. [2021], TCN Franceschi et al. [2019], and Flowformer Huang et al. [2022]. Compared to classical methods, RNN-based, Transformer-based, MLP-based, and other LLM-based approaches, DECA demonstrates consistently superior performance. It achieves an average accuracy improvement of 2.4%, 2.7%, and 2.8% over S2IP-LLM, Time-LLM, and GPT4TS, respectively. This demonstrates the versatility of DECA, showing that its framework can be effectively applied beyond time-series forecasting tasks to achieve excellent performance in other tasks as well.
@@
XGBoost Rocket LSTNet LSSL Trans. Re. In. Pyra. iTrans. Station. FED. ETS. Flow. DLinear LightTS. GPT4TS Time-LLM S2IP-LLM DECA

### v2

@@ C.2 Short-term Forecasting full results
We present the comprehensive short-term forecasting results in Table 11. Under different frequency settings, FSCA consistently outperforms most baseline models.
@@
width = 1colspec = Q[0.2em,c]Q[2em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c]Q[5.5em,c], colsep = 0.1pt, rowsep = 1.4pt, cells = c, cell11 = c=20.09, cell2,5,8,11,141 = r=3, hline1 = 1-140.15em, hline2,5,8,11,14 = 1-140.05em, hline17 = 1-140.15em, vline3,4,5,6,7,8,9,10,11,12,13,14 = 3,6,9,12,15, vline3,4,5,6,7,8,9,10,11,12,13,14 = 2,5,8,11,14abovepos = -1.2, vline3,4,5,6,7,8,9,10,11,12,13,14 = 4,7,10,13,16belowpos = -1.2, vline3,4,5,6,7,8,9,10,11,12,13,14 = 1abovepos = -1.2, belowpos = -1.2, Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST N-HiTS N-BEATS TimesNet FEDformer Stationary
@@ C.3 Few-shot Forecasting full results
Table 12 presents the detailed results of the few-shot forecasting trained on 5% data across different prediction lengths. Except for DLinear’s strong performance on the ETTm1 dataset, FSCA demonstrates significant enhancements in most scenarios across various datasets, outperforming the second-best model, S2IP-LLM, by 6.7%.
@@
colspec = Q[1.5em,c]Q[2.5em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c], colsep = 2.2pt, rowsep = 1.4pt, cells = c, cell11,3,5,7,9,11,13,15,17,19,21,23 = c=2, cell21 = c=2, cell3,8,13,181 = r=5, hline1 = 1-240.15em, hline2,3,8,13,18 = 1-240.05em, hline23 = 1-240.15em, vline2,3,5,7,9,11,13,15,17,19,21,23 = 4-6,9-11,14-16,19-21, vline2,3,5,7,9,11,13,15,17,19,21,23 = 3,8,13,18abovepos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 7,12,17,22belowpos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 1,2abovepos = -1.2, belowpos = -1.2, Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer
@@
Table 13 presents the detailed results of the few-shot forecasting task using 10% of the training data. FSCA demonstrates superior performance, outperforming the baseline across nearly all settings, with an average MSE reduction of 7.8% over the next best model, GPT4TS.
@@
colspec = Q[1.5em,c]Q[2.5em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c]Q[2em,c], colsep = 2.2pt, rowsep = 1.4pt, cells = c, cell11,3,5,7,9,11,13,15,17,19,21,23 = c=2, cell21 = c=2, cell3,8,13,181 = r=5, hline1 = 1-240.15em, hline2,3,8,13,18 = 1-240.05em, hline23 = 1-240.15em, vline2,3,5,7,9,11,13,15,17,19,21,23 = 4-6,9-11,14-16,19-21, vline2,3,5,7,9,11,13,15,17,19,21,23 = 3,8,13,18abovepos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 7,12,17,22belowpos = -1.2, vline2,3,5,7,9,11,13,15,17,19,21,23 = 1,2abovepos = -1.2, belowpos = -1.2, Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet FEDformer Stationary ETSformer
@@ C.4 Zero-shot Forecasting full results
Table 14 provides detailed results of the zero-shot forecasting task across different prediction lengths. FSCA consistently achieves optimal performance across all settings, with a significant margin. Compared to the second-best method, PatchTST, it shows an average improvement of 13.3%. Additionally, FSCA achieves MSE reductions of 18.3%, 17.7%, and 24.3% over other LLM-based methods S2IP-LLM, Time-LLM, and GPT4TS, respectively. This improvement is attributed to FSCA’s successful activation of LLMs’ structured understanding of TS data, while the design of the Dual-Scale Context-Alignment GNNs ensures that LLMs grasp the logical relationships in demonstration examples prompt.
@@
colsep = 2.2pt, rowsep = 1.4pt, cells = c, colspec = Q[5em,c]Q[1.5em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c]Q[2.9em,c], cell11,3,5,7,9,11,13,15,17 = c=2, cell21 = c=2, cell3,8,13,18,23,28,33,381 = r=5, hline1 = 1-240.15em, hline2,3,8,13,18,23,28,33,38 = 1-240.05em, hline43 = 1-240.15em, vline2,3,5,7,9,11,13,15,17 = 4-6,9-11,14-16,19-21,24-26,29-31,34-36,39-41, vline2,3,5,7,9,11,13,15,17 = 3,8,13,18,23,28,33,38abovepos = -1.2, vline2,3,5,7,9,11,13,15,17 = 7,12,17,22,27,32,37,42belowpos = -1.2, vline2,3,5,7,9,11,13,15,17 = 1,2abovepos = -1.2, belowpos = -1.2, Methods & FSCA S2IP-LLM Time-LLM GPT4TS iTransformer DLinear PatchTST TimesNet
@@ C.5 Classification Full Results
Table 15 presents the results on 10 multivariate UEA classification datasets. In the time series classification task, we have augmented the baselines with the following methods: XGBoost [Chen and Guestrin, 2016], Rocket [Dempster et al., 2020], LSTNet [Lai et al., 2018b], LSSL [Gu et al., 2021], LightTS [Zhang et al., 2022], Pyraformer [Liu et al., 2021], TCN [Franceschi et al., 2019], and Flowformer [Huang et al., 2022]. Compared to classical methods, RNN-based, Transformer-based, MLP-based, and other LLM-based approaches, FSCA demonstrates consistently superior performance. It achieves an average accuracy improvement of 2.4%, 2.7%, and 2.8% over S2IP-LLM, Time-LLM, and GPT4TS, respectively. This demonstrates the versatility of FSCA, showing that its framework can be effectively applied beyond time-series forecasting tasks to achieve excellent performance in other tasks as well.
@@
XGBoost Rocket LSTNet LSSL Trans. Re. In. Pyra. iTrans. Station. FED. ETS. Flow. DLinear LightTS. GPT4TS Time-LLM S2IP-LLM FSCA
@@ Appendix D VISUALIZATION
Figure 3 presents visualization examples of FSCA prediction results on the ETTh1, ETTm1, Electric, and Traffic datasets with an input length of 512 and a prediction length of 96. It can be observed that FSCA achieves good predictive performance across various datasets.